OpenStudy (anonymous):

Use the 5-step problem solving, show all work. Math Magicians charge a $45 registration fee plus $30 per hour for tutoring. Math Master charges $25 registration fee plus $35 per hour for tutoring. How many hours of tutoring would it take to make Math Magicians cheaper than Math Master?

OpenStudy (anonymous):

Magician = 45 + 30 * h Master = 25 + 35 * h 45 + 30 h < 25 + 35 h 45 < 25 + 5 h 20 < 5 h 4 < h